A ball is kicked such that it leaves the ground at an upwards angle of 35Β° above the horizontal with an initial velocity of 17 ms

How far away does the ball hit the ground?


1
Expert's answer
2022-03-15T11:15:02-0400

Range

"R=\frac{u^2sin(2\theta)}{2g}"

"R=\frac{17^2sin(70)}{9.8}=27.71m"

"R=" 27.71m